Determine the local minimum and local maximum values of the function f defined by


f(x) = 3-5x³ +5x⁴ -x^5

1
Expert's answer
2022-03-20T06:39:48-0400

"f'\left( x \right) =-15x^2+20x^3-5x^4=-5x^2\left( x-1 \right) \left( x-3 \right) \\f'\left( x \right) =0\Rightarrow x\in \left\{ 0,1,3 \right\}"



Then we see that x=1 is local minimum, x=3 is local maximum